(1 - -*)dx + (2xy + + + or? y? 2-)dy = 0 y J
Answer #1

Given equation is. 2y + + re ។ 7/3) dret (zny + en + tayin dy zo - 13 ) duhet 7 / 2) dy (dividing by both sides.) comparing w(N-a 11 + ay -اع 2y + vly) = S(N- sinca) dy s (27 + // ) dy y² + eny = vly) - so the solution of given equation es. Ulm, y) +